OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5?

OpenStudy (anonymous):

use the point/slope equation to find the equation of the line. y-y1=m(x-x1), make y the subject, and the slope of the line m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

make y the subject of "y-4=-5(x-3)", and you will have the equation of your line

OpenStudy (anonymous):

ok

Elsa213 (elsa213):

plug the value of the point & slope to get b 4=-8*3/4 +b 4=-6+b b=10 the equation = y = 3x/4+10